OpenStudy (quinton1011):

Multiply. Express your answer in lowest terms. -1/5 * (-3/10)

OpenStudy (quinton1011):

I'll give a medal

OpenStudy (quinton1011):

:( some one please help

OpenStudy (quinton1011):

OOPS! its -5/6 not -1/5

OpenStudy (anonymous):

\[\frac{ -1 }{ 5 } \times (\frac{ -3 }{ 10 })\]

OpenStudy (anonymous):

1/4

OpenStudy (quinton1011):

I'm confused

OpenStudy (anonymous):